OpenStudy (anonymous):

integration the following:

OpenStudy (anonymous):

\[\int\limits_{1}^{2} x^2 \div \sqrt{4-x^2}dx\]

sam (.sam.):

Use \(x=asin \theta\)

OpenStudy (anonymous):

sorry it'snot the right one :D hold a secound i will get the question foryou

OpenStudy (anonymous):

\[\int\limits_{1}^{2} 1\div \sqrt{1+\sqrt{x}}\]

OpenStudy (anonymous):

i use x= tantheta

sam (.sam.):

\[\Large \int\limits \frac{1}{\sqrt{1+\sqrt{x}}} \, dx?\]

sam (.sam.):

Just u-sub 2 times

sam (.sam.):

\[u=\sqrt{x} \\ \\ du=\frac{1}{2\sqrt x}\] \[2 \int\limits \frac{u}{\sqrt{u+1}} \, du\]

sam (.sam.):

Then s=u+1 ds=du \[2 \int\limits \frac{s-1}{\sqrt{s}} \, ds\] split denominator and integrate

OpenStudy (anonymous):

snx

OpenStudy (anonymous):

\[\int\limits_{2}^{1} tanx \div sinx -tanx\]

sam (.sam.):

\[\int\limits (\frac{t}{s}-t)dx?\]

OpenStudy (anonymous):

no :t/(s-t)

sam (.sam.):

\[\int\limits \frac{ \tan x}{ \sin x - \tan x } \, dx\]

sam (.sam.):

Hmm do you know weierstrass substitution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i mult by cotane then by conjegate

sam (.sam.):

Did you get something like \[\LARGE \int\limits -\frac{4 u}{\left(u^2-1\right) \left(u^2+1\right) \left(\frac{2 u}{u^2-1}+\frac{2 u}{u^2+1}\right)} \, du\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i get c+1(1/2-sin^2x)^-1

sam (.sam.):

weierstrass substitution is let \[u=\tan(\frac{x}{2})\] \[du=\frac{1}{2}\sec^2(\frac{x}{2})dx\] ----------------------------------- then let \[ \sin x=\frac{2 u}{u^2+1} ,~~~~~ \cos x=\frac{1-u^2}{u^2+1},~~~~ dx=\frac{2 du}{u^2+1}\]

OpenStudy (anonymous):

okay snx

OpenStudy (anonymous):

sam wouldn't be easier for the first question to let u = 1+\[\sqrt{x}\]

OpenStudy (anonymous):

to make x =(u-1)^2

OpenStudy (anonymous):

so easily get integration of u then reutrn bk '

sam (.sam.):

you can do that as well